OpenStudy (anonymous):

What is the sum of the arithmetic sequence 140, 134, 128…, if there are 33 terms? 1,364 1,408 1,452 1,496

OpenStudy (anonymous):

The 1st term is 140. The common difference is -6 The number of terms is 33. So the sum is (33/2)(2*140 + (33-1)(-6)) = (33/2)(280 - 196) = 1386